PDE
(6) Using the method of characteristics, solve the PDE диди Yör + au = 2 subject to u(x,0) = sin(x).
Answer #1

* Gliven yurtuyn, Uxio) = sind.. h = dy - dx=y dy integration Apply both sides San=Sydy a = 4 + dy = du ndy do integration Ap